Why does the voltage of a battery go up sharply when it gets close to being fully charged
and down sharply when it is close to fully discharged?
Answer #1

as we know all batteries have internal resistance.which is a function of the chemical components and number of cells inside the battery.and as we know the internal resistance of the battery is fixed and the lumped resistance may vary according to the parameter changes inside the battery.and this lumped resistance is in series with voltage source whose voltage is equal to open circuit terminal voltage of battery.if we call this Voc then the internal resistance is Rin then Ioc=Voc/Rin

therefore from ohms law V0c/Rt=Voc/Rin+Rl

Vdrop=Voc-Vl

Vdrop=

from the avalible equation the volatage is completely depens upon the the resistance across the battery .so as the battery is going dry the internal resistance goes high and the voltage comes down and when it is charging resistance decreses and the volage increase